A 0.30 kg metal ball at a velocity of 3.0 m/s hits another metal ball with a mass of 0.20 kg which is at rest. What is the velocity of the 0.20-kg metal ball if the velocity of the 0.30-kg metal ball became 2.0 m/s after the collision? Use: p before the collision (m1v1 + m2v2) = p after the collision (m1’v1’ + m2’v2’)


1
Expert's answer
2022-05-05T13:57:35-0400

The law of conservation of momentum says

"m_1v_1+m_2v_2=m_1v_1'+m_2v_2'""v_2=0"

Hence

"v_2'=\frac{m_1}{m_2}(v_1-v_1')"

"v_2'=\frac{0.30}{0.20}(3.0-2.0)=1.5\:\rm m/s"